### v1

# LLM with Relation Classifier for Document-Level Relation Extraction

Abstract
Large language models (LLMs) create a new paradigm for natural language processing. Despite their advancement, LLM-based methods still lag behind traditional approaches in document-level relation extraction (DocRE), a critical task for understanding complex entity relations. This paper investigates the causes of this performance gap, identifying the dispersion of attention by LLMs due to entity pairs without relations as a primary factor. We then introduce a novel classifier-LLM approach to DocRE. The proposed approach begins with a classifier specifically designed to select entity pair candidates exhibiting potential relations and thereby feeds them to LLM for the final relation extraction. This method ensures that during inference, the LLM’s focus is directed primarily at entity pairs with relations. Experiments on DocRE benchmarks reveal that our method significantly outperforms recent LLM-based DocRE models and achieves competitive performance with several leading traditional DocRE models.
LLM with Relation Classifier for Document-Level Relation Extraction
Xingzuo Li1111, Kehai Chen1111, Yunfei Long2222 and Min Zhang1111 $1$School of Computer Science and Technology, Harbin Institute of Technology, Shenzhen, China $2$School of Computer Science and Electronic Engineering, University of Essex, Colchester, United Kingdom 24S051028@stu.hit.edu.cn, chenkehai@hit.edu.cn, yl20051@essex.ac.uk, zhangmin2021@hit.edu.cn
1 Introduction
Document-level Relation Extraction (DocRE) stands as a critical yet formidable undertaking within the expansive landscape of Natural Language Processing (NLP). It delves into the intricate world of identifying and extracting semantic relationships that exist between entity pairs across an entire document, rather than confining the analysis to individual sentences. This task poses significant challenges due to the complex interplay of context, coreference resolution, and the long-range dependencies that are often embedded within the textual corpus. Previous DocRE models have relied on sophisticated neural network architectures to mimic human-like processes of comprehending and reasoning, including attention networks Zhang and Cheng (2022); Tan et al. (2022a), GNNs Li et al. (2020); Christopoulou et al. (2019), and have attained state-of-the-art performance Ma et al. (2023).
Figure 1: Statistics on the number of entity pairs in DocRED and Re-DocRED training set. NA: entity pairs that do not express any relation. Rel: entity pairs expressing relations.
Recently, researchers Sun et al. (2024) have delved into leveraging novel strategies to further boost the capabilities of Large Language Models (LLMs) in tackling DocRE. However, the LLM-based method still lags behind traditional approaches. Our meticulous observation highlights a pivotal issue: the DocRE task necessitates the construction of an exhaustive set of potential entity pairs, which we refer to as candidate space. Within this vast landscape of entity pairs, only a minute fraction actually express relations, creating a significant imbalance within the candidate space. As shown in Figure 1, our preliminary investigations, conducted on two widely adopted benchmark datasets, DocRED Yao et al. (2019) and Re-DocRED Tan et al. (2022b), underscored this phenomenon. The overwhelming majority of entity pairs labeled as NA (no relation), dominate the candidate space, potentially distracting LLMs from the minority that do encode meaningful relations. This skew in the data distribution is consequently recognized as a key contributor to the performance shortcomings exhibited by LLMs on DocRE.
Based on this finding, this paper introduces an innovative method LMRC, aimed at bridging the performance gap between the LLM-based DocRE methods and their conventional counterparts. Specifically, LMRC conceptualizes DocRE as a workflow comprising two key stages: Relation Candidate Proposal (RCP) and Relation Classification (RC). In the RCP stage, a pre-processing classifier is employed, which harnesses attention mechanism to filter out NA entity pairs that do not harbour any relation, significantly narrowing down the scope of candidate space. Subsequently, the RC stage leverages the power of LLMs to perform multi-classification within this refined candidate space, ensuring a more focused and accurate relation extraction process. Experimental results on the DocRED and Re-DocRED benchmarks showed that the proposed LMRC gains significant improvement over other LLM-based DocRE methods, suggesting its viability as a strategy for future DocRE.
Our core contributions are as follows:
  • Through a detailed analysis of the two benchmark datasets, DocRED and Re-DocRED, we identified that the main reason why LLM-based methods struggle to perform well on the DocRE task lies in the extreme imbalance of the candidate space.
  • We proposed LMRC approach, which first utilizes a pre-classifier to select potential relation-expressing entity pairs, and then leverages the reasoning capabilities of LLMs to complete relation classification.
  • Empirical experiments conducted on DocRED and Re-DocRED datasets have validated the feasibility and effectiveness of our method.
2 Related Works
Standard NLP methods for identifying relations in documents have included CNN and LSTMYao et al. (2019); Gu et al. (2017); Mandya et al. (2018), which are built upon the foundations laid by early deep neural networks. Jia et al. (2019) advocated for an entity-centric, multiscale approach, enabling a singular relation prediction for each entity tuple. More recently, approaches for document-level relation extraction have evolved around two main perspectives: Graph-based and Transformer-based. The concept of document graph was initially introduced by Quirk and Poon (2016). In this paradigm, words are represented as nodes, while linguistic relations (e.g., adjacency, syntactic dependencies) are captured as edges. Graph-based methods Zeng et al. (2020); Zhang et al. (2020) particularly excels in handling long-distant dependencies that often proved crucial in comprehending complex relations. In addition to modifying the graph structure, Xu et al. (2022) proposed a sentence importance estimation and focusing framework, which encourages models to focus on evidence sentences. In early Transformer-based methods, Verga et al. (2018) introduced Bi-affine Relation Attention Network to identify the relations between entities. Han and Wang (2020) utilized a document-level entity mask method augmented with type information to provide a more richer contextual understanding of the entities. Zhou et al. (2021) and Tan et al. (2022a) further leveraged attention mechanism to enhance entity representations, improving the accuracy of relation classification. Recently, Ma et al. (2023) extended previous research by integrating knowledge distillation strategy and achieved SOTA performance.
Relation Extraction with LLMs   LLMs have demonstrated immense potential across multiple domains Jiang et al. (2024); Jing et al. (2024). Several recently proposed RE approaches have advocated the utilization of LLMs to output string encodings, thereby addressing the task in an interpretable manner. Paolini et al. (2021) introduced a framework that reframes many structured prediction tasks, including relation extraction, as a sequence-to-sequence problem. Wan et al. (2023) pushed the boundaries by prompting GPT-3 for relation extraction on datasets such as TACRED Zhang et al. (2017) and SciERC Luan et al. (2018), achieving SOTA performance. Wadhwa et al. (2023) extended this line of work by leveraging few-shot in-context learning within LLMs for sentence-level relation extraction.
Previously, the application of LLMs to document-level relation extraction has been limited. Ozyurt et al. (2023) introduced a method for in-context few-shot relation extraction leveraging pre-trained language models, providing a viable paradigm for tackling DocRE with LLMs. Sun et al. (2024) attempted to employ chain-of-retrieval and denoising strategy to steer LLMs towards accurately identifying relations. During their research on GPT-3’s contextual learning abilities for document-level biomedical information extraction tasks, Gutierrez et al. (2022) uncovered a general issue in in-context learning when dealing with the NA category. Inspire by this discovery, we first focus on selecting potential relation-expressing entity pairs and then utilize the reasoning capability of LLMs to complete relation classification. As far as we know, our work is among the first to apply LLMs for document-level relation extraction.
In Context Learning   In-context learning Brown et al. (2020) has emerged as a competitor against supervised baselines across a wide range of tasks. This approach circumvents the computational burden of weight updates by leveraging a handful of pertinent examples presented to the model. Recent works on in-context learning have mainly focused on optimizing prompt retrieval Liu et al. (2022); Rubin et al. (2022) and prompt ordering Lu et al. (2022). Typically, Wan et al. (2023) and Wadhwa et al. (2023) have evaluated the in-context learning performance of GPT-3 on sentence-level RE tasks. We conduct our experiment in 5.6 following their works.
3 Preliminary
3.1 Problem Fomulation
Given a document $D$ that includes a set of sentences $X_{D}={\{x_{i}\}}_{i}^{k}$ and a set of entities $E_{D}={\{e_{i}\}}_{i}^{n}$, document-level relation extraction aims to predict a subset of relations from $R\cup\{NA\}$ for the entity pairs ${(e_{s},e_{o})}_{s,o=1,...,n;s\neq o}$, where $R$ represents a predefined set of relations, $e_{s}$ and $e_{o}$ respectively denote the subject entity and the object entity, and NA indicates that there is no relation between the entities. A entity $e_{i}$ can appear multiple times within a document through its mentions $M_{i}={\{m_{j}^{i}\}}^{N_{i}}_{j=1}$, where $m_{j}^{i}$ represents the $j^{th}$ mention and $N_{i}$ denotes the number of mentions. During test time, the model is required to predict labels for all entity pairs.
3.2 Pre-Experiments
To investigate the reasons behind the under-performance of current LLMs in DocRE, we fine-tune LLaMA2-13B-Chat for this task and report the outcomes and key insights derived from applying this approach to the DocRED and Re-DocRED datasets.
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
Your task is to determine whether there are relations between the entity pairs based on the information in the text. If there exists relations, select relations for the entity pairs from the relation set; if there is no relation, return None.
The format of the input entity pair is ‘(head entity| -| tail entity)’.
Your output format is ‘(head entity| relation/None| tail entity)’.
### Relation set:
{predefined relation set}
### Text:
{text}
### {number of entity pairs} Entity pairs:
{entity pairs}
### Response:
Table 1: Prompt for document-level relation extraction
Fine-tuning LLaMA2 In order to construct effective prompts for this task, we employ the following instructional guideline: Your task is to determine whether there are relations between the entity pairs based on the information in the text. If there exist relations, select relations for the entity pairs from the relation set; if there is no relation, return None.. The instruction is accompanied by an input that includes a predefined relation set, the text corresponding to the document, and the entity pairs that need to be classified. To prevent ambiguities and optimize token usage in our process, we use None as a substitute for NA and require the model to explicitly label entity pairs that do not contain relations as None. The complete prompt format is provided in Table 1.
Each document in DocRE task involves a considerable number of tokens, frequently exceeding the maximum limit supported by typical processing unit. To tackle this challenge, we devise a method to conduct relation extraction for each document $D$ via $\frac{n\times(n-1)}{k}$ inputs, where $n$ denotes the number of entities within document $D$, and the variable $k$ represents the capacity limit for entity pairs per input instance. We then integrate all entity pairs into the inputs adhering to the prescribed scheme to perform LoRA Hu et al. (2022) fine-tuning and testing [^1] Entities in the triplets returned by LLaMA2-13B-Chat are aligned to the dataset using thefuzz, and the relations generated not in the predefined relation set are considered incorrect. on LLaMA2-13B-Chat.
Results Statistics of DocRED and Re-DocRED are shown in Table 2. NA entity pairs, which do not exhibit relations, constitute a significant proportion in both datasets, leading to an imbalance in the candidate space. Further to the empirical observations by Lilong et al. (2024), our study extends the analysis to examine the model’s outputs from a distribution perspective, supported by experiments, aiming to uncover the fundamental reasons behind the observed under-performance. As demonstrated in Table 3, the number of triples generated by LLaMA2-13B-Chat is markedly less than that annotated in the dataset. This discrepancy suggests that LLMs such as LLaMA2 tend to label relation-expressing entity pairs as NA, leading to the poor outcomes.
Description DocRED Re-DocRED
Dev Test Dev Test
Candidate Space 395,572 392,158 193,232 198,670
# NA Entity Pairs 384,949 - 179,870 185,043
# Relation Entity Pairs 10,623 - 13,362 13,627
# Annotated Triples 12,275 - 17,284 17,448
Table 2: Statistics on DocRED and Re-DocRED
Metrics DocRED Re-DocRED
Dev Test Dev Test
$Precision$ 69.00 - 84.88 83.94
$Recall$ 27.43 - 38.06 38.14
$F_{1}$ 39.25 38.66 52.56 52.45
Ign $F_{1}$ 38.62 38.09 52.29 52.15
# Extracted Triples 4,925 4,932 7,787 7,979
Table 3: Results of preliminary experiment.
4 Proposed Method: LMRC
To prevent LLMs from prioritizing NA entity pairs, the proposed LMRC methodology, as shown in Figure 2, initially uses traditional neural networks for Relation Candidate Proposal to identify relation-expressing entity pairs. Then, LLMs rely on these proposals for Relation Classification.
4.1 Relation Candidate Proposal
In this stage, we build a simple model to conduct a binary classification task, with the outcome being entity pairs expressing relations. As prior works Tan et al. (2022a); Ma et al. (2023) have shown that contextual information is indispensable for the relation extraction task, our model adapts localized context pooling from Zhou et al. (2021).
Entity Representation   Following the entity marker technique Zhang et al. (2017); Shi and Lin (2019), a special token "*" is inserted at the start and end position of each entity mention. Then, tokens $T=\{t_{i}\}_{i=1}^{l}$ within document $D$ are encoded by a Transformer-basedVaswani et al. (2017) pretrained language model (PLM) to generate contextualized embeddings $\mathbf{H}$ along with their attentions $\mathbf{A}$:
$\mathbf{H},\mathbf{A}=PLM(T),$ (1)
where $\mathbf{H}\in\mathbb{R}^{l\times d}$, $\mathbf{A}\in\mathbb{R}^{H\times l\times l}$, $d$ is the hidden dimension of the PLM and $H$ is the number of attention heads. We take the embedding of "*" at the start of mentions as their embeddings. The entity embedding $h_{i}\in\mathbb{R}^{d}$ for each entity $e_{i}$ with mentions $M_{i}=\{m_{j}^{i}\}_{j=1}^{N_{i}}$ is computed by logsumexp pooling Jia et al. (2019):
$h_{i}=log\sum_{j=1}^{N_{i}}exp(h_{m_{j}^{i}}).$ (2)
Localized Context Representation   For each entity $e_{i}$, we aggregate the attention output for its mentions by mean pooling $A_{i}=\sum_{j=1}^{N_{i}}(a_{m_{j}^{i}})$, where $a_{m_{j}^{i}}\in\mathbb{R}^{H\times l}$ is the attention weight at the position of mention $m_{j}^{i}$ from the last layer. Then given an entity pair $(e_{s},e_{o})$, its localized context embedding $c^{(s,o)}\in\mathbb{R}^{d}$ can be obtained by:
$q^{(s,o)}$ $=\sum_{i=1}^{H}(A_{s}^{i}\circ A_{o}^{i}),$ (3)
$c^{(s,o)}$ $=\mathbf{H}^{\top}q^{(s,o)},$ (4)
where $q^{(s,o)}\in\mathbb{R}^{l}$ is the mean-pooled attention weight for entity pair $(e_{s},e_{o})$ and $\mathbf{H}$ is the contextualized embedding in Eq.(1).
Binary Classification   To predict whether entity pair $(e_{s},e_{o})$ expresses relations, we first generate context-enhanced entity representations:
$z_{s}^{(s,o)}=tanh(\mathbf{W}_{s}h_{s}+\mathbf{W}_{c}c^{(s,o)}),$ (5)
where $\mathbf{W}_{s},\mathbf{W}_{c}\in\mathbb{R}^{d\times d}$ are trainable parameters. We obtain the object representation $z_{o}^{(s,o)}$ in the same manner. Then, a bilinear classifier is applied on the representations to compute the probability:
$P(\textit{NA}|e_{s},e_{o})=\sigma(z_{s}^{(s,o)\top}\mathbf{W}z_{o}^{(s,o)}+b),$ (6)
where $\mathbf{W}\in\mathbb{R}^{d\times d}$ is a trainable parameter matrix, $\sigma$ is the sigmoid function, $P(\textit{NA}|e_{s},e_{o})$ is the probability that entity pair $(e_{s},e_{o})$ does not express any relation. We choose Binary Cross Entropy as our loss function:
$\mathcal{L}=\sum I(\textit{NA})logP+(1-I(\textit{NA}))log(1-P),$ (7)
where $I$ represents Indicator Function, and $I(\textit{NA})$ takes the value of 1 when there is no relation between entities. $P$ denotes $P(\textit{NA}|e_{s},e_{o})$.
4.2 Relation Classification
Figure 2: Illustration of LMRC. Relation Candidate Proposal(RCP) leverages localized context pooling Zhou et al. (2021) in the construction of a pre-processing classifier, focusing on selecting entity pairs that contain relations. Relation Classification(RC) takes the results from the previous stage to create prompt that guides fine-tuned LLaMA2 to accomplish multi-classification task.
After identifying potential relation-expressing entity pairs in the RCP stage, we employ the method described in Section 3, leveraging LLMs to undertake supervised fine-tuning for relation classification. This approach involves a subtle modification to the previous prompting scheme, where we streamline the process by eliminating the NA category and change certain expressions. These adjustments are intended to further sharpen the model’s concentration on the classification task, ensuring a more precise and targeted performance.
Additionally, the optimization achieved through the RCP stage yields a substantial reduction in the number of inputs per document, transitioning from a complexity of $\frac{n\times(n-1)}{k}$ to $\frac{n^{\prime}}{k}$, where $n^{\prime}$ denotes the number of entity pairs identified in the RCP stage. This reduction, driven by the elimination of NA instances, not only enhances computational efficiency but also shortens the inference time of LLMs. Details can be found in Table 4.
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
This is a relation classification task. we will provide entity pairs that require relation classification. Your task is to select relations for each entity pair from the given relation set based on the information in the text. There may be multiple relations between an entity pair.
The format of the input entity pair is ‘(head entity| -| tail entity)’.
Your output format is ‘(head entity| relation| tail entity)’.
### Relation set:
{predefined relation set}
### Text:
{text}
### {number of entity pairs} Entity pairs:
{entity pairs}
### Response:
Table 4: Prompt for relation classification
5 Experiments
5.1 Experimental Settings
Datasets   We conduct experiments on DocRED Yao et al. (2019) and Re-DocRED Tan et al. (2022b), two large-scale crowd-sourced benchmark datasets tailored for document-level RE.
DocRED, rooted in Wikipedia and Wikidata, stands as a milestone in DocRE. It provides a comprehensive annotation schema, encompassing entity mentions and types, relational facts, as well as the corresponding supporting evidence. With 97 distinct target relations and an average of approximately 26 entities per document, DocRED offers a rich and challenging landscape for researchers. Notably, over 40.7% of the relational facts within DocRED require multi-sentence reasoning. Although DocRED is a widely recognized benchmark, the annotations of the dataset remain incomplete. Tan et al. (2022b) thus proposed Re-DocRED, a more reliable benchmark for DocRE that revises DocRED to mitigate the false negative issue within it.
Configuration   In the RCP stage, RoBERTalarge Liu et al. (2019) is chosen as the foundational PLM encoder. We implement early stopping based on the $F_{1}$ score obtained from the development set. We adopt AdamW Loshchilov and Hutter (2019) as the optimizer and apply a linear warmup for the learning rate at the first 6% steps. Main parameters are provided in Table 5. We use development set to manually tune the optimal parameters for the RCP stage based on the F1 score. The value of parameters we finally adopted are in bold.
Parameters DocRED Re-DocRED
batch size 4 4
# Epoch 20, 30, 40 30, 40
lr for encoder {5, 3, 1}e-5 {3, 1}e-5
lr for classifier 1e-4 1e-4
max gradient norm 1.0 1.0
Table 5: Settings for the RCP stage.
Parameters Pre RC stage
DocRED Re-DocRED DocRED Re-DocRED
batch size 4 4 4 4
# Epoch 2 2 8 8
learning rate 1e-4 1e-4 1e-4 1e-4
warmup steps 200 200 100 100
lora r 8 8 8 8
lora alpha 16 16 16 16
Table 6: Settings for LoRA fine-tuning. (Pre stands for preliminary experiment)
In the RC stage, we fine-tune LLMs with the RC-specific prompt, which is detailed in 4.2, using LoRA. Details regarding parameters are provided in Table 6. We train and evaluate the pre-classifier on a single NVIDIA A100 GPU during the RCP stage, and fine-tune LLMs on 4 NVIDIA A100 GPUs during the RC stage.
Evaluation   Consistent with other SOTA methods, we employ the standard evaluation metrics: $F_{1}$ and Ign $F_{1}$. Ign $F_{1}$ is calculated by excluding triplets that are already present in the training set from both the development and test sets.
5.2 Main Results
Method Dev Test
Ign $F_{1}$ $F_{1}$ Ign $F_{1}$ $F_{1}$
BERT-based
HIN-BERTbaseTang et al. (2020) 54.29 56.31 53.70 55.60
CorefBERTbaseYe et al. (2020) 55.32 57.51 54.54 56.96
CorefRoBERTalargeYe et al. (2020) 57.35 59.43 57.90 60.25
HeterGSAN+reconstructionXu et al. (2021b) 58.13 60.18 57.12 59.45
SSAN-RoBERTalargeXu et al. (2021a) 60.25 62.08 59.47 61.42
KD-RoBERTalargeTan et al. (2022a) 65.27 67.12 65.24 67.28
DREEAM-RoBERTalargeMa et al. (2023) 65.52 67.41 65.47 67.53
LLM-based
CoRSun et al. (2024) - 38.4 ± 10.6 - 38.5 ± 9.1
GenRDKSun et al. (2024) - 42.5 ± 10.6 - 41.5 ± 8.7
Our Methods
LoRA FT LLaMA2-7B-Chat 33.95 34.32 33.99 34.34
LoRA FT LLaMA2-13B-Chat 38.62 39.25 38.09 38.66
LMRC-LLaMA2-7B-Chat 52.40 54.10 52.81 54.73
LMRC-LLaMA2-13B-Chat 58.16 59.97 58.49 60.52
(a) Results on the development and test set of DocRED.
Method Ign $F_{1}$ $F_{1}$
BERT-based
KD-RoBERTalargeTan et al. (2022a) 77.60 78.28
DREEAMMa et al. (2023) 79.66 80.73
LLM-based
CoRSun et al. (2024) - 37.1 ± 9.2
GenRDKSun et al. (2024) - 41.3 ± 8.9
AutoRELilong et al. (2024) - 51.91
Our Methods
LoRA FT LLaMA2-7B-Chat 52.74 53.02
LoRA FT LLaMA2-13B-Chat 52.15 52.45
LMRC-LLaMA2-7B-Chat 72.33 72.92
LMRC-LLaMA2-13B-Chat 74.08 74.63
(b) Results on the test set of Re-DocRED
[…]
We compare our LMRC with pretrained BERT-based and LLM-based methods on both datasets. BERT-based methods, known for achieving state-of-the-art (SOTA) performance, utilize BERT family pretrained models as encoders. Recently introduced LLM-based methods employ fine-tuning, in-context learning with CoT Wei et al. (2022), or retrieval augmented generation (RAG, Lewis et al. (2020)) to enhance the performance of LLMs on relation extraction.
Overall Performance   As demonstrated in Table 7(a), the direct fine-tuning approach applied to LLaMA2 and other LLM-based methods expose suboptimal performance and inefficiencies in processing capabilities, underscoring the inherent challenges faced when leveraging LLMs for DocRE. Our results also corroborate the findings of Lilong et al. (2024). However, following the implementation of our proposed task division strategy, we observe a substantial improvement, marked by a significant boost in the $F_{1}$ score for LLaMA2 at both 7B and 13B scales. Table 7(b) presents a comparative analysis of LMRC’s performance against existing methods evaluated on the Re-DocRED test set. We observe that LMRC surpasses other LLM-based methods. Moreover, LMRC narrows the performance gap with the current SOTA method, DREEAM, positioning it as a promising paradigm for future DocRE.
Method Intra Inter
[…]
LoRA FT LLaMA2-13B-Chat 45.43 31.67
LMRC + LLaMA2-13B-Chat 65.88 52.66
[…]
Performance of Intra-/Inter-Extraction   We also report Intra-$F_{1}$/Inter-$F_{1}$, which respectively evaluate the model’s capability in capturing intra- and inter-sentence relations. LSR Nan et al. (2020) and GAIN Zeng et al. (2020), both graph-centric methodologies that leverages graph neural networks for inference, serve as valuable benchmarks. As illustrated in Table 8, LMRC not only outperforms the baseline achieved by directly fine-tuning LLaMA2 in both Intra- and Inter-$F_{1}$, but also maintains competitiveness with advanced graph-based models like GAIN-BERTbase. This demonstrates LMRC’s effectiveness, particularly in scenarios that require complex relational parsing from document-level context.
5.3 Fine-grained Analysis for Each Relation
To comprehensively assess the model’s sensitivity to different categories within the dataset, we undertake a fine-grained statistical analysis of the $F_{1}$ score for each relation type in the DocRED development set, comparing both the baseline and LMRC. Figure 3 visually showcases the enhancement effects in a more intuitive manner, highlighting the gains achieved by LMRC.
Our meticulous data analysis leads to the following conclusions: (1) LMRC enhanced the extraction performance for the majority of relations at both 7B and 13B scales. (2) A discernible positive relationship emerges between the frequency of occurrence of relations in the training set and their ultimate extraction performance. High-frequency relations, such as P569:"date of birth", achieved a remarkable $F_{1}$ score exceeding 90, showcasing the model’s proficiency in handling prevalent patterns.
However, it is noteworthy that despite these success, LMRC encounters challenges in extracting instances pertained to certain relations (e.g., P1336:"territory claimed by", P39:"position held"). We hypothesize that this limitation stems primarily from the scarcity of these relations in the training data, which restricts the model’s capacity to sufficiently learn the relevant features and nuances. Consequently, the model struggles to generalize effectively for these less frequently encountered relation types.
Figure 3: $F_{1}$ scores per relation type in the DocRED development set results (darker = better). White color means that no correct predictions were made for this relation. The relations are arranged in descending order by the number of triples.
5.4 Ablation Studies
After showing the effectiveness of our proposed LMRC, we further explore the roles of RCP and RC stages on the DocRED development set. To validate the RCP stage, we fine-tune LLaMA2-13B-Chat to replace the original pre-classifier for binary classification. Moving on to the RC stage, we introduce a more simplified setup where we feed the model with entity pairs annotated in the ground truth. By masking relation tags, we challenge the model to classify these pairs into the predefined relation set. This process involves task-specific fine-tuning of LLaMA2-13B-Chat, specifically tailored for relation classification.
As shown in Table 9, substituting our pre-classifier with the fine-tuned LLaMA2-13B-Chat leads to a significant decrease in the $F_{1}$ score. We contend that the intricate multi-relation nature as well as the discrete distribution of relations in the DocRE task pose distinct challenges for LLMs, highlighting the crucial role of the pre-classifier in ensuring robust performance. Furthermore, the ablation study conducted on the RC stage underscores the proficiency of the fine-tuned LLaMA2-13B-Chat in performing relation classification. This achievement provides a solid foundation for future advancements, demonstrating the potential of LLMs excelling in complex relation extraction tasks.
Settings $F_{1}$ of RCP Ign $F_{1}$ $F_{1}$
RCP stage
LMRC 64.64 58.16 59.97
w/o pre-classifier w LLM 31.30 23.22 24.59
RC stage
relation classification - 86.09 86.75
Table 9: Ablation studies evaluated on DocRED dev set.
5.5 Out-of-Domain Relation Studies
In the aforementioned evaluation, we conservatively categorize all relations generated by LLaMA2 that do not align with the predefined relation set as erroneous outcomes1. However, previous research Wadhwa et al. (2023) has indicated that evaluating LLM-based models should not entirely rely on exact matches to targets. This is because LLMs can generate outputs that are semantically proximate to the target relations but may not satisfy the stringent exact-matching requirement. For instance, while the phrase "works at" in the model’s output resembles "work for" in the target, strict evaluation criteria would count it as a misclassification.
To gain a comprehensive understanding of this phenomenon, we delve deeper into the out-of-domain relations generated by LLaMA2-13B-Chat. We leverage SBERT [^2] https://www.sbert.net Reimers and Gurevych (2019), a well recognized approach for semantic similarity, to align out-of-domain relations to the predefined relation set $R$. This process first computes the cosine similarity between each out-of-domain relation and the relations within $R$. For each $r_{i}$ in the out-of-domain set, we choose the relation in $R$ that exhibits the highest similarity score $s_{\text{max}}^{i}$ as the optimal match. Intuitively, some out-of-domain relations may not carry meaningful information due to the hallucination issue of LLMs, and blindly aligning them to $R$ is not appropriate. Therefore, we introduce a heuristic threshold $\theta$. Only when $s_{\text{max}}^{i}\geq\theta$, does the alignment processed; otherwise, the triplets containing $r_{i}$ are discarded. After alignment, we recalculate the $F_{1}$ score of our methods on the DocRED dev set.
(a) LoRA FT
(b) LMRC
Figure 4: The impact of threshold $\theta$ for cosine similarity on the F1 score. Both methods are conducted on the DocRED dev set.
Figure 4 illustrates that LLaMA2 directly fine-tuned on DocRED attains its peak $F_{1}$ score when threshold $\theta$ is set approximately at 0.55. However, despite incorporating semantic alignment, the out-of-domain relations generated by LLaMA2 fine-tuned within the LMRC framework remain incorrect. We analyze these relations in depth and conclude that they mainly suffer from the following two problems:
  • The model outputs "-", which appears in the entity pair input format, as a relation.
  • While some of the generated out-of-domain relations could ostensibly be mapped to similar relations within the set $R$, the classification result is incorrect.
Additionally, it is noteworthy that both methods generated a relatively low number of out-of-domain relations, 74 and 72, respectively. This limited occurrence helps explain why adjustments to the threshold $\theta$ has little impact on the overall $F_{1}$ score.
5.6 Performance of GPT-4-turbo
To ensure a thorough assessment, we expand our investigation to encompass a broader array of models, notably incorporating GPT-3-turbo-16k and GPT-4-turbo into our study. This inclusion was made to validate the adaptability and robustness of our approach to DocRE across diverse model architectures.
We first leveraged a 3-shot learning format for GPT mdoels to directly tackle DocRE, utilizing examples meticulously curated from the human-annotated DocRED dataset. These examples were carefully chosen to include a diverse range of entity pairs, capturing both relation-expressing and non-relation entity pairs, thereby mirroring the complexity and variability inherent in real-world documents. Then, we adopted the LMRC paradigm. In this setting, GPT models took over the role of LLaMA2 in the RC stage, augmented by the addition of 3-shot learning format to enrich contextual understanding.
Methods Ign $F_{1}$ $F_{1}$ Intra $F_{1}$ Inter $F_{1}$
GPT-3.5-turbo-16k + 3-shot 6.57 6.97 11.78 4.24
GPT-4-turbo + 3-shot 9.85 10.71 14.52 7.43
LMRC + GPT-3.5-turbo-16k 27.27 28.39 38.52 17.19
LMRC + GPT-4-turbo 34.84 36.20 44.08 27.30
LMRC + LLaMA2-13B-Chat (FT) 58.16 59.97 65.88 52.66
Table 10: Performance of GPT models on sampled documents from the DocRED dev set. FT: fine-tuning.
Considering budget constraints, our experimental analysis was conducted on a randomly selected subset of 100 documents from the DocRED development set. The preliminary results are shown in Table 10. Before implementing LMRC, the extraction performances of GPT models were quite low, while their overall F1 scores were significantly improved after integrating LMRC strategy. This reflects the inherent challenges posed by DocRE and also demonstrates the effectiveness of our proposed method on various models. In addition, even GPT-4-turbo failed to surpass the fine-tuned LLaMA2-13B-Chat through 3-shot in-context learning, indicating the necessity of fine-tuning LLMs on DocRE tasks.
6 Conclusion
In this work, we investigate the underlying reasons for LLM’s limited effectiveness in document-level relation extraction and introduce a new approach, Large Language Models with Relation Classifier (LMRC), for DocRE. Our method comprises two main stages: relation candidate proposal and relation classification. Through experiments conducted on DocRED and Re-DocRED, we demonstrate the effectiveness of our proposed LMRC approach. The results further reveal that LMRC holds strong competitive advantages over other existing LLM-based methods. Our innovative model establishes a new standard, indicating its potential as a viable framework for future DocRE research.